# Expanding the Chaos: Neural Operator for Stochastic (Partial) Differential Equations

### v1

@@ 6 Conclusion
In this work, we presented a family of NO models for solving SDEs and SPDEs that are built on classical Wiener–chaos expansions of their solutions. Our approach projects the driving noise onto orthonormal Wick–Hermite features and learns deterministic propagator functions parameterised by neural operators, so that full trajectories can be reconstructed from noise in a single forward pass. Building on existing WCE results, we made the associated propagator systems for multi-dimensional SDEs and semilinear SPDEs explicit and used them to motivate our architectures, and we demonstrated competitive accuracy and broad applicability across classical SPDE benchmarks, diffusion one-step sampling, graph interpolation, coupled Heston extrapolation, parameter estimation, and manifold SDEs, while retaining the one-shot evaluation advantage of neural operators.
General Structure of Appendix
The general structure of the appendix can be divided into Theoretical part and Empirical part.
  • Theoretical part: From Appendix B to E, we conduct a thorough analysis on both WCE theory and our proposed models.
  • Empirical Part: In appendix F, we show details of our experiments, including data generation, model architecture, training and evaluation, as well as some additional results and analysis.
Appendices Contents
@@ Appendix A Related Works
Wiener chaos expansion (WCE), also known as polynomial chaos expansion, is a classical technique in stochastic analysis that represents random processes via an orthogonal polynomial basis. Originating from Norbert Wiener’s homogeneous chaos in 1938 and the Cameron–Martin expansion in 1947 [wiener1938homogeneous, cameron1947orthogonal]. In the initial stage, WCE was first applied to the field of stochastic finite element analysis [ghanem2003stochastic] and extended to the theory of probability distributions, known as Wiener–Askey polynomial chaos [xiu2002wiener]. WCE allows one to expand the solution of SDE or SPDE in terms of polynomial functionals with deterministic coefficients, known as propagators. Spectral Galerkin methods based on WCE have been successfully applied to propagate uncertainty in complex systems, e.g., efficiently solving stochastic Burgers and Navier–Stokes equations driven by random forcing [hou2006wiener]. To handle strong nonlinearity or long-time integration, advanced variants, such as multi-element polynomial chaos, decompose the random input domain and build local expansions in each element, thereby improving convergence for problems with discontinuities or large perturbations [wan2006multi]. In recent years, there is a growing interest in combining WCE with machine learning techniques. For instance, neural chaos methods replace the fixed orthogonal polynomials with trainable neural network basis functions to learn an optimal expansion from data [sharma2025polynomial]. Likewise, WCE has been employed as a data-driven operator-learning approach in scientific machine learning, achieving competitive accuracy in learning solution maps of PDEs while providing built-in uncertainty quantification [sharma2024physics]. These developments demonstrate how WCE has evolved from a purely theoretical construct into a versatile computational tool for both numerical uncertainty quantification and modern ML-based modeling.
@@
Classical approaches for solving SDEs and SPDEs build on established numerical discretization methods such as finite difference schemes, finite element methods, and spectral Galerkin approximations [platen2010numerical, orszag1971numerical]. Monte Carlo simulation is another cornerstone, often employed to handle stochastic inputs or to estimate statistical properties of the solution [jentzen2009numerical]. These techniques have been extensively validated on canonical models: for example, finite difference methods reliably capture solutions and sensitivities in financial SDE benchmarks like the Heston stochastic volatility model [milstein2005numerical], while spectral Galerkin (e.g. polynomial chaos) methods and fine-grid CFD solvers have been applied to SPDEs such as turbulent Navier–Stokes equations with random forcing [abraham2018spectral]. These learning approaches show promise (often alleviating the curse of dimensionality or enabling inverse analyses), but the well-established numerical schemes above remain the standard reference for accuracy and robustness in solving stochastic (partial) differential equations.
@@
generalize neural networks to learn mappings between infinite-dimensional function spaces, enabling discretization-invariant (resolution-generalizable) solution of parametric PDEs [kovachki2023neural]. Foundational frameworks like DeepONet and the Fourier Neural Operator (FNO) demonstrated this paradigm by accurately learning nonlinear solution operators for diverse PDE families [lu2019deeponet, li2020fourier]. These models learn an entire family of equations rather than a single instance, with FNO achieving zero-shot super-resolution across mesh resolutions. Building on these advances, recent works have extended neural operators to stochastic domains. For SPDEs, specialized architectures incorporate the effects of random forcing into the operator learning process: for example, a Neural SPDE model can parameterize solution operators depending simultaneously on the initial condition and a realization of the noise [salvi2022neural], and the Neural Operator with Regularity Structure (NORS) leverages Hairer’s regularity structure theory to handle the low regularity of SPDE solutions [gong2023deep, hu2022neural]. In the SDE setting, SDE-ONet augments the DeepONet architecture with polynomial chaos expansions to efficiently learn solution operators for stochastic differential equations [eigel2024functional]. These developments illustrate the growing capability of neural operators to serve as efficient surrogates for both deterministic and stochastic systems, while preserving their core advantages of infinite-dimensional approximation and mesh-independent generalization. In recent years, machine learning-based solvers have also emerged – e.g., physics-informed neural networks and GAN-driven models that learn solution mappings – to tackle high-dimensional or data-driven SDE/SPDE problems [li2022dynamic].
@@ Appendix B Formulation details and Proofs for Section 2
In this section, we provide a detailed formulation of the concepts in Section 2 and some conclusions regarding the uniqueness of their solutions. Several results are standard and included only for completeness; for these, we give brief proof sketches and refer the readers to introductions of SDEs in oksendal2003stochastic and SPDEs in holden1996stochastic, liu2015stochastic.
@@ B.1 A Thorough Formulation of SPDE and SDE
To support the Appendix, we present full SDE and SPDE formulations below, supplementing the SDE (SDE) and SPDE (SPDE). Similar, to align with previous works [neufeld2024solving, huschto2014solving, eigel2024functional], we adopt the following notations for self-completeness.
@@
We now recall the mild solution of SPDE. One can call an $\mathbb{F}$-predictable [da2014stochastic] process $X:[0,T]\times\Omega\to\mathcal{H}$ a mild solution of (SPDE) if $\mathbb{P}[\int_{0}^{T}\|X_{t}\|_{\mathcal{H}}^{2}dt<\infty]=1$ and for every $t\in[0,T]$ it holds that
@@
which is known as Itô’s SDE, with $W_{t}$ as a $m$-dimensional Brownian motion with covariance $Q\in\mathbb{R}^{m\times m}$, i.e., $\mathbb{E}[(W_{t_{2}}-W_{t_{1}})(W_{t_{2}}-W_{t_{1}})^{\top}]=(t_{2}-t_{1})Q$, and $F(t,\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d}$, $B(t,\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are known as the drift and diffusion term of the SDE, respectively. Without loss of generality, in this work, we only consider the standard Brownian motion, i.e., $Q=I$, and our conclusions can be smoothly extended to the general case. Then the strong solution of (SDE) is
@@ B.2 Existence and Uniqueness of SPDE, SDE Solutions
Proposition 1 (SPDE Solution Existence and Uniqueness [neufeld2024solving, da2014stochastic]).
Let $p\in[1,\infty)$ and assume that Assumption 1 holds. Then, the (SPDE) has a unique $\mathbb{F}$-predictable mild solution $X\colon[0,T]\times\Omega\to\mathcal{H}$. Furthermore, there exists a constant $C_{p,T}$, depending only on $p$, $C_{F,B}$, $T$, and $C_{S}\coloneqq\sup_{t\in[0,T]}\|S_{t}\|_{L(\mathcal{H};\mathcal{H})}$, such that the following moment bound holds:
@@
The proof of this proposition is a cornerstone result in the theory of stochastic evolution equations and can be found in detail in standard literature, for instance, in da2014stochastic. The complete argument relies on applying the Banach fixed-point theorem to an appropriate operator on a space of stochastic processes, combined with maximal inequalities for stochastic convolutions. Thus we omit the proof here. ∎
@@
Let Assumption 2 hold. Then, the (SDE) admits a unique, $\mathbb{F}$-adapted, and continuous strong solution $X\colon[0,T]\times\Omega\to\mathbb{R}^{d}$. Furthermore, for any $p\in[1,\infty)$, there exists a constant $C_{p,T}$, depending only on $p$, $T$, and the constants $L$ and $K$ from Assumption 2, such that the solution’s moments are bounded:
@@
This is a classical result in the theory of stochastic differential equations, with a detailed proof available oksendal2003stochastic. The proof for existence and uniqueness is typically established by constructing a sequence of approximations via Picard iteration and showing that this sequence converges to a limit that is the unique solution. This is achieved by proving that the corresponding integral operator is a contraction mapping on a suitable Banach space of stochastic processes. ∎
Appendix C Proofs for Conclusions in Section 3
@@
This is a classical result (see, e.g., da2014stochastic or neufeld2024solving); we briefly recall the main idea. For a fixed component $W^{(i)}$, the indicator $\mathbf{1}_{[0,t]}$ admits the $L^{2}([0,T])$ expansion $\mathbf{1}_{[0,t]}(s)=\sum_{j\geq 1}G_{j}(t)\,e_{j}(s)$, which yields $W_{t}^{(i)}=\int_{0}^{T}\mathbf{1}_{[0,t]}(s)\,dW_{s}^{(i)}=\sum_{j\geq 1}G_{j}(t)\,\xi_{ij}$ by linearity and the Itô isometry. The truncated processes $\widehat{W}^{(n,i)}_{t}=\sum_{j=1}^{n}G_{j}(t)\,\xi_{ij}$ converge in $L^{2}(\Omega)$ for each fixed $t$, and applying a maximal inequality (e.g., Doob or BDG) together with the summability assumption $\sum_{j}\|G_{j}\|_{C([0,T])}^{2}<\infty$ gives convergence in $L^{2}(\Omega;C([0,T]))$. Almost sure uniform convergence then follows by a standard Borel–Cantelli argument; see the above references for details.
@@
This is the natural extension of Lemma 1 to a $Q$-Brownian motion on a Hilbert space and follows standard arguments based on the Karhunen–Loève expansion (see, e.g., [da2014stochastic, Section 3.3] and [lototsky2017stochastic, Section 2]). Writing
@@
C.2 Proofs of the conclusions in Section 3.2
In this section, we briefly recall standard properties of Wick polynomials that are used in Section 3.2; full details can be found in classical references on Wiener–Itô chaos such as [lototsky2017stochastic, luo2006wiener, neufeld2024solving].
@@
This is a standard identity; see, for example, lototsky2017stochastic or luo2006wiener.
@@
Each factor is computed using Proposition 3, yielding $\mathbb{E}[h_{\alpha_{mj}}(\xi_{mj})\,h_{\beta_{mj}}(\xi_{mj})]=\alpha_{mj}!\,\delta_{\alpha_{mj},\beta_{mj}}$. If $\alpha\neq\beta$, there exists at least one $(m,j)$ with $\alpha_{mj}\neq\beta_{mj}$, so the product vanishes. If $\alpha=\beta$, then all Kronecker deltas are one and we obtain $\mathbb{E}[\xi_{\alpha}^{2}]=1$ by the definition of $\alpha!$. This proves orthonormality; see, e.g., lototsky2017stochastic for a textbook derivation. ∎
@@
This is the classical Wiener–Itô/Cameron–Martin theorem; see, for example, [lototsky2017stochastic, Theorem 1.1], [luo2006wiener, Chapter 2], or [neufeld2024solving, Section 2]. We do not reproduce the proof here.
@@
The conclusion for the 1-dimensional SDE is established in [eigel2024functional](Theorem 2.2), we slightly extend the Theorem to $d$-dimensional case. We first recall our theorem presented on the main page.
@@
Let $X_{t}\in\mathbb{R}^{d}$ be the strong solution to the $d$-dimensional SDE (SDE) driven by an $m$-dimensional Brownian motion $W_{t}$. The solution’s WCE is given by
@@
The integral form of the $j$-th component of the SDE (SDE) is:
@@
After presenting the propagator system for SDEs, we now turn to SPDE propagators. Similar chaos-based formulations of SPDEs can be found in the Wiener–chaos literature; see, for example, [lototsky2006stochastic, lototsky2006wiener, kalpinelli2011wiener, neufeld2024solving]. Our SPDE propagator system (Theorem 2) specialises these results to the semilinear SPDE (SPDE) and writes the associated deterministic PDEs for the Wiener–chaos coefficients in the compact form (6), which can be used directly for neural-operator parameterisation; see also the discussion in Remark 2.13 of [neufeld2024solving]. Similarly, we recall our theorem in SPDEs.
@@
The evolution of the propagator fields $u_{\alpha}(t,x)$ for (SPDE) is governed by a coupled system of deterministic PDEs
@@
The derivation follows the classical Wiener–chaos approach for SPDEs (see, e.g., lototsky2006wiener, lototsky2006stochastic, kalpinelli2011wiener, neufeld2024solving), adapted to the semilinear SPDE (SPDE). Starting from the mild formulation
@@
Consider the SPDE (SPDE). Assume the drift term $F(t,X_{t})$ is a polynomial in $X_{t}$, expressible as a finite sum of Wick powers:
@@
In this work we use the Wick–Hermite chaos basis, which is standard in the Wiener–chaos/SPDE literature and aligns naturally with the algebraic structure of Wick products and with existing propagator and error analyses [lototsky2017stochastic, luo2006wiener, neufeld2024solving]. We note that there are alternative polynomial bases for approximating Brownian motion and Gaussian processes, such as the optimal bases studied by Foster et al. [foster2020optimal]. These bases can be advantageous from a pure approximation point of view, but it is not yet clear whether they preserve the same simple chaos decomposition and Wick–product algebra needed to derive explicit propagator systems and source terms $\mathscr{F}_{\alpha}$. Exploring such alternative bases within the SDENO/F-SPDENO framework, and understanding the trade-offs between structural convenience and approximation optimality, is an interesting direction for future work.

### v2

@@ 6 Conclusion
In this work, we presented a family of NO models for solving SDEs and SPDEs that are built on classical Wiener–chaos expansions of their solutions. Our approach projects the driving noise onto orthonormal Wick–Hermite features and learns deterministic propagator functions parameterised by neural operators, so that full trajectories can be reconstructed from noise in a single forward pass. Building on existing WCE results, we made the associated propagator systems for multi-dimensional SDEs and semilinear SPDEs explicit and used them to motivate our architectures, and we demonstrated competitive accuracy and broad applicability across classical SPDE benchmarks, diffusion one-step sampling, graph interpolation, coupled Heston extrapolation, and manifold SDEs, while retaining the one-shot evaluation advantage of neural operators.
@@ Appendix A Related Works
Wiener chaos expansion (WCE), also known as polynomial chaos expansion, is a classical technique in stochastic analysis that represents random processes via an orthogonal polynomial basis. Originating from Norbert Wiener’s homogeneous chaos in 1938 and the Cameron–Martin expansion in 1947 (Wiener, 1938; Cameron and Martin, 1947). In the initial stage, WCE was first applied to the field of stochastic finite element analysis (Ghanem and Spanos, 2003) and extended to the theory of probability distributions, known as Wiener–Askey polynomial chaos (Xiu and Karniadakis, 2002). WCE allows one to expand the solution of SDE or SPDE in terms of polynomial functionals with deterministic coefficients, known as propagators. Spectral Galerkin methods based on WCE have been successfully applied to propagate uncertainty in complex systems, e.g., efficiently solving stochastic Burgers and Navier–Stokes equations driven by random forcing (Hou et al., 2006). To handle strong nonlinearity or long-time integration, advanced variants, such as multi-element polynomial chaos, decompose the random input domain and build local expansions in each element, thereby improving convergence for problems with discontinuities or large perturbations (Wan and Karniadakis, 2006). In recent years, there is a growing interest in combining WCE with machine learning techniques. For instance, neural chaos methods replace the fixed orthogonal polynomials with trainable neural network basis functions to learn an optimal expansion from data (Sharma et al., 2025). Likewise, WCE has been employed as a data-driven operator-learning approach in scientific machine learning, achieving competitive accuracy in learning solution maps of PDEs while providing built-in uncertainty quantification (Sharma et al., 2024). These developments demonstrate how WCE has evolved from a purely theoretical construct into a versatile computational tool for both numerical uncertainty quantification and modern ML-based modeling.
@@
Classical approaches for solving SDEs and SPDEs build on established numerical discretization methods such as finite difference schemes, finite element methods, and spectral Galerkin approximations (Platen and Bruti-Liberati, 2010; Orszag, 1971). Monte Carlo simulation is another cornerstone, often employed to handle stochastic inputs or to estimate statistical properties of the solution (Jentzen and Kloeden, 2009). These techniques have been extensively validated on canonical models: for example, finite difference methods reliably capture solutions and sensitivities in financial SDE benchmarks like the Heston stochastic volatility model (Milstein et al., 2005), while spectral Galerkin (e.g. polynomial chaos) methods and fine-grid CFD solvers have been applied to SPDEs such as turbulent Navier–Stokes equations with random forcing (Abraham et al., 2018). These learning approaches show promise (often alleviating the curse of dimensionality or enabling inverse analyses), but the well-established numerical schemes above remain the standard reference for accuracy and robustness in solving stochastic (partial) differential equations.
@@
generalize neural networks to learn mappings between infinite-dimensional function spaces, enabling discretization-invariant (resolution-generalizable) solution of parametric PDEs (Kovachki et al., 2023). Foundational frameworks like DeepONet and the Fourier Neural Operator (FNO) demonstrated this paradigm by accurately learning nonlinear solution operators for diverse PDE families (Lu et al., 2019; Li et al., 2020). These models learn an entire family of equations rather than a single instance, with FNO achieving zero-shot super-resolution across mesh resolutions. Building on these advances, recent works have extended neural operators to stochastic domains. For SPDEs, specialized architectures incorporate the effects of random forcing into the operator learning process: for example, a Neural SPDE model can parameterize solution operators depending simultaneously on the initial condition and a realization of the noise (Salvi et al., 2022), and the Neural Operator with Regularity Structure (NORS) leverages Hairer’s regularity structure theory to handle the low regularity of SPDE solutions (Gong et al., 2023; Hu et al., 2022). In the SDE setting, SDE-ONet augments the DeepONet architecture with polynomial chaos expansions to efficiently learn solution operators for stochastic differential equations (Eigel and Miranda, 2024). These developments illustrate the growing capability of neural operators to serve as efficient surrogates for both deterministic and stochastic systems, while preserving their core advantages of infinite-dimensional approximation and mesh-independent generalization. In recent years, machine learning-based solvers have also emerged – e.g., physics-informed neural networks and GAN-driven models that learn solution mappings – to tackle high-dimensional or data-driven SDE/SPDE problems (Li and Feng, 2022).
@@ Appendix B Formulation details and Proofs for Section 2
In this section, we provide a detailed formulation of the concepts in Section 2 and some conclusions regarding the uniqueness of their solutions. Several results are standard and included only for completeness; for these, we give brief proof sketches and refer the readers to introductions of SDEs in (Øksendal and Øksendal, 2003) and SPDEs in (Holden et al., 1996; Liu and Röckner, 2015).
@@ B.1 A Thorough Formulation of SPDE and SDE
To support the Appendix, we present full SDE and SPDE formulations below, supplementing the SDE (2) and SPDE (1). Similar, to align with previous works (Neufeld and Schmocker, 2024; Huschto and Sager, 2014; Eigel and Miranda, 2024), we adopt the following notations for self-completeness.
@@
We now recall the mild solution of SPDE. One can call an $\mathbb{F}$-predictable (Da Prato and Zabczyk, 2014) process $X:[0,T]\times\Omega\to\mathcal{H}$ a mild solution of (1) if $\mathbb{P}[\int_{0}^{T}\|X_{t}\|_{\mathcal{H}}^{2}dt<\infty]=1$ and for every $t\in[0,T]$ it holds that
@@
which is known as Itô’s SDE, with $W_{t}$ as a $m$-dimensional Brownian motion with covariance $Q\in\mathbb{R}^{m\times m}$, i.e., $\mathbb{E}[(W_{t_{2}}-W_{t_{1}})(W_{t_{2}}-W_{t_{1}})^{\top}]=(t_{2}-t_{1})Q$, and $F(t,\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d}$, $B(t,\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are known as the drift and diffusion term of the SDE, respectively. Without loss of generality, in this work, we only consider the standard Brownian motion, i.e., $Q=I$, and our conclusions can be smoothly extended to the general case. Then the strong solution of (2) is
@@ B.2 Existence and Uniqueness of SPDE, SDE Solutions
Proposition 1 (SPDE Solution Existence and Uniqueness (Neufeld and Schmocker, 2024; Da Prato and Zabczyk, 2014)).
Let $p\in[1,\infty)$ and assume that Assumption 1 holds. Then, the (1) has a unique $\mathbb{F}$-predictable mild solution $X\colon[0,T]\times\Omega\to\mathcal{H}$. Furthermore, there exists a constant $C_{p,T}$, depending only on $p$, $C_{F,B}$, $T$, and $C_{S}\coloneqq\sup_{t\in[0,T]}\|S_{t}\|_{L(\mathcal{H};\mathcal{H})}$, such that the following moment bound holds:
@@
The proof of this proposition is a cornerstone result in the theory of stochastic evolution equations and can be found in detail in standard literature, for instance, in (Da Prato and Zabczyk, 2014). The complete argument relies on applying the Banach fixed-point theorem to an appropriate operator on a space of stochastic processes, combined with maximal inequalities for stochastic convolutions. Thus we omit the proof here. ∎
@@
Let Assumption 2 hold. Then, the (2) admits a unique, $\mathbb{F}$-adapted, and continuous strong solution $X\colon[0,T]\times\Omega\to\mathbb{R}^{d}$. Furthermore, for any $p\in[1,\infty)$, there exists a constant $C_{p,T}$, depending only on $p$, $T$, and the constants $L$ and $K$ from Assumption 2, such that the solution’s moments are bounded:
@@
This is a classical result in the theory of stochastic differential equations, with a detailed proof available (Øksendal and Øksendal, 2003). The proof for existence and uniqueness is typically established by constructing a sequence of approximations via Picard iteration and showing that this sequence converges to a limit that is the unique solution. This is achieved by proving that the corresponding integral operator is a contraction mapping on a suitable Banach space of stochastic processes. ∎
Appendix C Proofs and Clarifications for Conclusions in Section 3
@@
This is a classical result (see, e.g., Da Prato and Zabczyk (2014, Proposition 4.3) or Neufeld and Schmocker (2024, Section 2)); we briefly recall the main idea. For a fixed component $W^{(i)}$, the indicator $\mathbf{1}_{[0,t]}$ admits the $L^{2}([0,T])$ expansion $\mathbf{1}_{[0,t]}(s)=\sum_{j\geq 1}G_{j}(t)\,e_{j}(s)$, which yields $W_{t}^{(i)}=\int_{0}^{T}\mathbf{1}_{[0,t]}(s)\,dW_{s}^{(i)}=\sum_{j\geq 1}G_{j}(t)\,\xi_{ij}$ by linearity and the Itô isometry. The truncated processes $\widehat{W}^{(n,i)}_{t}=\sum_{j=1}^{n}G_{j}(t)\,\xi_{ij}$ converge in $L^{2}(\Omega)$ for each fixed $t$, and applying a maximal inequality (e.g., Doob or BDG) together with the summability assumption $\sum_{j}\|G_{j}\|_{C([0,T])}^{2}<\infty$ gives convergence in $L^{2}(\Omega;C([0,T]))$. Almost sure uniform convergence then follows by a standard Borel–Cantelli argument; see the above references for details.
@@
This is the natural extension of Lemma 1 to a Q-Brownian motion on a Hilbert space and follows standard arguments based on the Karhunen–Loève expansion (see, e.g., (Da Prato and Zabczyk, 2014, Section 3.3) and (Lototsky et al., 2017, Section 2)). Writing
@@
C.2 Proofs and Clarifications of the conclusions in Section 3.2
In this section, we briefly recall standard properties of Wick polynomials that are used in Section 3.2; full details can be found in classical references on Wiener–Itô chaos such as (Lototsky et al., 2017; Luo, 2006; Neufeld and Schmocker, 2024).
@@
This is a standard identity; see, for example, Lototsky et al. (2017, Chapter 1) or Luo (2006, Section 2).
@@
Each factor is computed using Proposition 3, yielding $\mathbb{E}[h_{\alpha_{mj}}(\xi_{mj})\,h_{\beta_{mj}}(\xi_{mj})]=\alpha_{mj}!\,\delta_{\alpha_{mj},\beta_{mj}}$. If $\alpha\neq\beta$, there exists at least one $(m,j)$ with $\alpha_{mj}\neq\beta_{mj}$, so the product vanishes. If $\alpha=\beta$, then all Kronecker deltas are one and we obtain $\mathbb{E}[\xi_{\alpha}^{2}]=1$ by the definition of $\alpha!$. This proves orthonormality; see, e.g., Lototsky et al. (2017, Section 2) for a textbook derivation. ∎
@@
This is the classical Wiener–Itô/Cameron–Martin theorem; see, for example, (Lototsky et al., 2017, Theorem 1.1), (Luo, 2006, Chapter 2), or (Neufeld and Schmocker, 2024, Section 2). We do not reproduce the proof here.
@@
The conclusion for the 1-dimensional SDE is established in (Eigel and Miranda, 2024)(Theorem 2.2), we slightly extend the Theorem to $d$-dimensional case. We first recall our theorem presented on the main page.
@@
Let $X_{t}\in\mathbb{R}^{d}$ be the strong solution to the $d$-dimensional SDE (2) driven by an $m$-dimensional Brownian motion $W_{t}$. The solution’s WCE is given by
@@
The integral form of the $j$-th component of the SDE (2) is:
@@
After presenting the propagator system for SDEs, we now turn to SPDE propagators. Similar chaos-based formulations of SPDEs can be found in the Wiener–chaos literature; see, for example, (Lototsky and Rozovskii, 2006a, b; Kalpinelli et al., 2011; Neufeld and Schmocker, 2024). Our SPDE propagator system (Theorem 2) specialises these results to the semilinear SPDE (1) and writes the associated deterministic PDEs for the Wiener–chaos coefficients in the compact form (6), which can be used directly for neural-operator parameterisation; see also the discussion in Remark 2.13 of (Neufeld and Schmocker, 2024).
Similarly, we recall our theorem in SPDEs.
@@
The evolution of the propagator fields $u_{\alpha}(t,x)$ for (1) is governed by a coupled system of deterministic PDEs
@@
The derivation follows the classical Wiener–chaos approach for SPDEs (see, e.g., Lototsky and Rozovskii (2006b, a); Kalpinelli et al. (2011); Neufeld and Schmocker (2024)), adapted to the semilinear SPDE (1). Starting from the mild formulation
@@
Consider the SPDE (1). Assume the drift term $F(t,X_{t})$ is a polynomial in $X_{t}$, expressible as a finite sum of Wick powers:
@@
In this work we use the Wick–Hermite chaos basis, which is standard in the Wiener–chaos/SPDE literature and aligns naturally with the algebraic structure of Wick products and with existing propagator and error analyses (Lototsky et al., 2017; Luo, 2006; Neufeld and Schmocker, 2024). We note that there are alternative polynomial bases for approximating Brownian motion and Gaussian processes, such as the optimal bases studied by Foster et al. (Foster et al., 2020). These bases can be advantageous from a pure approximation point of view, but it is not yet clear whether they preserve the same simple chaos decomposition and Wick–product algebra needed to derive explicit propagator systems and source terms $\mathscr{F}_{\alpha}$. Exploring such alternative bases within the SDENO/F-SPDENO framework, and understanding the trade-offs between structural convenience and approximation optimality, is an interesting direction for future work.